OpenStudy (anonymous):

ok so for my final review i need to know few things sooo. yea can someone help me with few things 1. calculating% composition by mass 2. converting particles to moles and moles to particles 3. converting mass to moles and moles to mass 4. converting mass to particles and particles to mass can u plz help me sum it down a lil plz and thank you

OpenStudy (aaronq):

1. to find the mass percent composition you use the molar masses mass % composition of A \(=\dfrac{(\#~of~atoms~of~A)*(Molar ~mass ~of ~A)}{Molar~mass~of ~compound}*100\%\) example: Find the mass percent composition of H in water (\(H_2O\)) percent composition (by mass) of H\(=\dfrac{2 atoms*1~/gmol}{18~g/mol}*100\%=11.11\%\)

OpenStudy (aaronq):

the formatting ended up a little weird on the top equation, i hope it's clear.

OpenStudy (aaronq):

for 2 through 4, you use these 2 equations: \(moles=\dfrac{mass}{Molar~mass}\) and \(moles=\dfrac{molecules~(or~atoms,ions, etc)}{Avogadro's `number}\) (mass needs to be in grams) so lets say you wanna know how many moles there is in 10 grams of water, \(moles=\dfrac{mass}{Molar~mass}=\dfrac{10~g}{18~g/mol}=0.555 ~moles\) now how many water molecules in 0.555 moles? \(moles=\dfrac{molecules~(or~atoms,ions, etc)}{Avogadro's `number}\) \(0.555=\dfrac{molecules}{6.022*10^{23}}\rightarrow molecules=0.555*(6.022*10^{23})=3.34*10^{23}~molecules\)

OpenStudy (sweetburger):

i wish my chem final had this stuff on it

OpenStudy (anonymous):

thank you so much aarong

OpenStudy (aaronq):

no problem